# Elastic tensor-derived properties of composition-dependent disordered refractory binary alloys

### v2

@@ 3.5 Nuclear-relaxation contributions to elastic properties
The method employed in computing the nuclear-relaxation contributions to the elastic tensor makes it possible to decompose it into atom-specific contributions, allowing more detailed modeling of the microscopic origins of elastic response in the binary alloys. The force response internal strain tensor $\Lambda_{n}^{\alpha}$ characterizes the force response of atom $\alpha$ along direction $n$ (in Voigt notation) to the internal strain, while the displacement response internal strain tensor $\Gamma_{n}^{\alpha}$ characterizes the displacement response of the atom $\alpha$ along direction $n$ to the internal strain. The product of these intermediate values summed over all directions for a given atom $\alpha$ gives the relaxation value of a given elastic constant or shear modulus $G$ of atom $\alpha$ under macroscopic strain (Eqs. (16) and (2.4)). To illustrate the resulting spatial variation in a relaxed crystal structure, the relaxation effects are displayed as contour plots in Figs. 6-7 for NbTa and WTa, and for MoNb, MoTa, MoW, and WTa in the Appendix. The supercell is represented as a set of 0 0 1delimited-⟨⟩001$\langle$ 0\,0\,1$\rangle$ 0 0 1 ⟩ planes. A supercell with unit cells stacked in $2\times 2\times 2$ fashion has 4 unique 0 0 1delimited-⟨⟩001$\langle$ 0\,0\,1$\rangle$ 0 0 1 ⟩ planes. These are used to construct 2D layers in order to represent the 3D periodic supercell. Each layer contains the atoms of their respective 0 0 1delimited-⟨⟩001$\langle$ 0\,0\,1$\rangle$ 0 0 1 ⟩ plane, and their first nearest neighbors (1NN) from the next plane. For example, layer 1 consists of atoms from the plane 1 and atoms from plane 2 (these are the 1NN of the atoms in plane 1) projected onto plane 1. Interpolation between atomic contributions is used to construct the spatially-dependent relaxation field [104] for each elastic tensor quantity, using a linear function. The color bar in the contour plot has a range varying from minimum (blue) to maximum (yellow) relaxation value.
@@
3.6 Computational efficiency of DFPT vs. stress-strain methods
In order to assess the computational efficiency of DFPT for computing the elastic tensor vs. stress-strain methods  [40], we computed elastic tensor for $\rm{Mo_{0.5}W_{0.5}}$ as an exemplar system, using a 16-atom SQS supercell. The DFT computational parameters and atomic configuration were the same as used in the DFPT calculations for this binary. In the stress-strain method, the elastic tensor is determined by computing the stress tensor for each applied strain using DFT calculation in which atomic positions are relaxed while the simulation cell shape and dimensions are fixed. Consequently, the elastic tensor is obtained from a linear fit of computed stresses with respect to imposed strains $\bm{s}_{ij}$ of -1%, -0.5%, 0.5%, 1% as given in Eq. (1). The resulting stress-strain-computed elastic tensor for $\rm{Mo_{0.5}W_{0.5}}$ differs by $<2\%$ from the relaxed-ion-computed elastic tensor.
The computation of rigid-ion and relaxed-ion elastic tensors using DFPT requires only one calculation once the relaxed structure had been obtained. Computation of the rigid-ion elastic tensor is faster than the computation of the relaxed structure; hence, the total time needed for computing the rigid-ion elastic tensor is less than the time required to perform two independent computations. By contrast, the stress strain-method requires 25 independent computations. In the case of the relaxed-ion method, the total time for the computation is greater than for the rigid-ion method, as the energy derivatives with respect to the atomic displacements for all atoms are needed. In our calculations for $\rm{Mo_{0.5}W_{0.5}}$, we found that the rigid-ion method was $\times2$ faster while the relaxed-ion method was 1.7×1.7$\times$1.7 × slower than the stress-strain method.
An alternative method for computing the elastic tensor utilizes a finite-difference approach to compute the necessary energy derivatives. Numerical validation of the DFPT compared to finite difference methods has been done previously [47, 52, 106]. In the finite difference approach, a minimum of 5 DFT computations are needed in order to compute each independent elastic constant using a 5-point derivative formula. Martin et al. [52] investigated the computational efficiency of DFPT for computing the elastic tensor of ${\rm Fe}_{16}{\rm C}$ with trigonal symmetry and seven independent elastic constants, using a PAW pseudopotential. They found that DFPT was 18.5×18.5$\times$18.5 × faster than the finite difference approach.
@@ 5 Conclusions
In this work, the elastic tensors of composition-dependent refractory binary alloys based on Mo, Nb, Ta, and W have been computed at a high level of quantum mechanical theory using density functional perturbation theory in conjunction with SQS supercell modeling of the composition-dependent disordered BCC structures. Both the rigid-ion and relaxed ion elastic tensors were computed. The Voigt-averaged elastic constants $\bar{C}_{11},\bar{C}_{12}$, and $\bar{C}_{22}$ derived from the elastic tensors were used to compute the polycrystalline elastic moduli $E$, $B$, and $G$. Relaxed-ion calculations for the elastic constants corresponding to the inclusion of nuclear relaxation effects under uniform strain are consistent with experiment. Differences between the rigid-ion and relaxed-ion values for the elastic constants were found to lie within the standard deviations of their Voigt averages for the 25, 50, and 75 at.alloy compositions, suggesting that nuclear relaxation contributions to the elastic tensor are relatively small. This made it possible to extend rigid-ion elastic tensor and derived elastic properties to additional alloy compositions with no loss in accuracy. The computed rigid-ion elastic constants are in excellent agreement with available experimental data, demonstrating the ability of our methodology to accurately predict key mechanical properties from first-principles theory. Derived mechanical properties of $B$, $G$ and $E$ showed a linear trend in properties across the compositions for same-group binary alloys and a nonlinear trend in different-group binary alloys. Our calculations showed elastic anisotropy behavior in all of the binary alloys except WTa.
In addition to the ability to neglect nuclear relaxation contributions in computing accurate mechanical properties for the considered alloys, we compared our approach to conventional stress-strain methods, and found that accurate results can be obtained with an order of magnitude reduction in computational cost. For the test system of $\rm{Mo_{0.25}Nb_{0.75}}$, the rigid-ion DFPT calculation of $B$ was in exact agreement with the results of conventional stress-strain approach. Even though the relaxation values are small, the decomposition of the elastic constants into electronic contributions and nuclear relaxation contributions helped to understand the effects of nuclear relaxations on the symmetry of the elastic tensor. Computation of relaxation values allowed to map the heterogeneity in elastic constants which can play a crucial role in understanding crack initial site in disordered alloys. The ability to perform reliable high-throughput calculations will make it possible to establish accurate property databases for microscopic modeling of the mechanical properties of compositionally- and chemically-complex alloys.

### v3

@@ 3.5 Nuclear-relaxation contributions to elastic properties
[…]
@@ 5 Conclusions
In this work, the elastic tensors of composition-dependent refractory binary alloys based on Mo, Nb, Ta, and W have been computed using density functional perturbation theory in conjunction with SQS supercell modeling of the composition-dependent disordered BCC structures. The use of the response-based DFPT methodology for computing the elastic tensor made it possible to compute both the rigid-ion and relaxed ion elastic tensors and nuclear-relaxation contributions, in addition to elastic tensor-derived properties. The Voigt-averaged elastic constants $\bar{C}_{11},\bar{C}_{12}$, and $\bar{C}_{44}$ derived from the elastic tensors were used to compute the polycrystalline elastic moduli $E$, $B$, and $G$. Relaxed-ion calculations for the elastic constants corresponding to the inclusion of nuclear relaxation effects under uniform strain are consistent with experiment. Differences between the rigid-ion and relaxed-ion values for the elastic constants were found to lie within the standard deviations of their Voigt averages for the 25, 50, and 75 at.alloy compositions, suggesting that nuclear relaxation contributions to the elastic tensor are relatively small. This made it possible to extend rigid-ion elastic tensor and derived elastic properties to additional alloy compositions with no loss in accuracy. The computed rigid-ion elastic constants are in excellent agreement with available experimental data, demonstrating the ability of our methodology to accurately predict key mechanical properties from first-principles theory. Derived mechanical properties of $B$, $G$ and $E$ showed a linear trend in properties across the compositions for same-group binary alloys and a nonlinear trend in different-group binary alloys. Our calculations showed elastic anisotropy behavior in all of the binary alloys except WTa.
The computed relaxation values were shown to be small and the decomposition of the elastic constants into electronic contributions and nuclear relaxation contributions helps to understand the effects of nuclear relaxations on the symmetry of the elastic tensor. The lattice-dependent nuclear relaxation fields provide detailed quantum mechanical information for potential use in data-driven predictive models. Computation of relaxation values made it possible to map the heterogeneity in elastic constants which can play a crucial role in understanding crack initial site in disordered alloys. This demonstrated ability to perform accurate, experimentally-validated DFPT calculations will enable advanced, microscopic modeling of the mechanical properties of compositionally- and chemically-complex alloys.
@@ Acknowledgements
Data Availability
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Appendix
The following figures provide additional computational results for binaries not discussed in the main text. Figures 912 illustrate the spatial dependence of $G$ (GPa), $E$ (GPa), and $\nu$ for MoTa, MoW, NbTa and WNb, as a function of composition. Figures 1316 illustrate the relaxation field of $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$ as a function of composition for MoNb, MoTa, MoW and WNb.
Figure 9: Spatial dependence of $G$ (GPa), $E$ (GPa), and $\nu$ for: (a)-(c)  $\rm{Mo_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}Ta_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
Figure 10: Spatial dependence of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{Mo_{0.25}W_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}W_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}W_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
Figure 11: Spatial dependence of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{Nb_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{Nb_{0.5}Ta_{0.5}}$;  and (g)-(h) $\rm{Nb_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
Figure 12: Spatial dependence of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{W_{0.25}Nb_{0.75}}$;  (d)-(f) $\rm{W_{0.5}Nb_{0.5}}$; and (g)-(h) $\rm{W_{0.75}Nb_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
Figure 13: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.5}Nb_{0.5}}$, and $\rm{Mo_{0.75}Nb_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 14: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, and $\rm{Mo_{0.75}Ta_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 15: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}W_{0.75}}$, $\rm{Mo_{0.5}W_{0.5}}$, and $\rm{Mo_{0.75}W_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 16: Relaxation field (values in GPa) constructed as described in text for $\rm{W_{0.25}Nb_{0.75}}$, $\rm{W_{0.5}Nb_{0.5}}$, and $\rm{W_{0.75}Nb_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.